# RobustBlack: Challenging Black-Box Adversarial Attacks on State-of-the-Art Defenses

### v1

@@ Appendix E Detailed Tabular Results
Table 12: Success rates of black-box attacks using vanilla surrogates against a vanilla target and robust targets.
Models GHOST LGV BASES TREMBA VMI-FGSM TI-FGSM ADMIX VNI-FGSM DI-FGSM UAP MI-FGSM
Standard ResNet50 49.36 $\pm$ 0.14 15.98 $\pm$ 1.13 81.08 $\pm$ 0.49 89.56 $\pm$ 0.09 61.22 $\pm$ 0.18 41.5 $\pm$ 0.35 64.75 $\pm$ 0.17 61.10 $\pm$ 0.13 65.95 $\pm$ 0.75 7.11 $\pm$ 0.25 51.62 $\pm$ 0.0
Salman ResNet18 1.13 $\pm$ 0.08 1.37 $\pm$ 0.12 2.36 $\pm$ 0.04 4.06 $\pm$ 0.06 1.29 $\pm$ 0.05 2.04 $\pm$ 0.06 1.88 $\pm$ 0.08 1.35 $\pm$ 0.06 1.51 $\pm$ 0.03 0.83 $\pm$ 0.03 0.98 $\pm$ 0.0
Liu ConvNeXt-L 0.48 $\pm$ 0.09 0.45 $\pm$ 0.01 1.16 $\pm$ 0.0 1.47 $\pm$ 0.01 0.7 $\pm$ 0.02 1.19 $\pm$ 0.1 0.99 $\pm$ 0.03 0.8 $\pm$ 0.04 0.92 $\pm$ 0.05 0.18 $\pm$ 0.04 0.6 $\pm$ 0.0
Liu ConvNeXtV2-L+Swin-L 1.7 $\pm$ 0.19 1.0 $\pm$ 0.03 4.77 $\pm$ 0.13 2.94 $\pm$ 0.12 2.51 $\pm$ 0.02 2.08 $\pm$ 0.04 2.52 $\pm$ 0.06 2.45 $\pm$ 0.05 2.79 $\pm$ 0.07 0.63 $\pm$ 0.03 2.0 $\pm$ 0.0
Liu Swin-L 0.53 $\pm$ 0.03 0.45 $\pm$ 0.01 1.26 $\pm$ 0.02 1.25 $\pm$ 0.01 0.74 $\pm$ 0.02 1.22 $\pm$ 0.05 0.93 $\pm$ 0.05 0.74 $\pm$ 0.04 0.88 $\pm$ 0.03 0.18 $\pm$ 0.0 0.51 $\pm$ 0.0
@@
Table 13: Success rates of black-box attacks using robust surrogates against a vanilla target and robust targets.
Models GHOST LGV BASES TREMBA VMI-FGSM TI-FGSM ADMIX VNI-FGSM DI-FGSM UAP MI-FGSM
Standard ResNet50 17.95 $\pm$ 0.36 7.79 $\pm$ 1.58 17.94 $\pm$ 0.00 67.54 $\pm$ 0.13 16.47 $\pm$ 0.02 8.48 $\pm$ 0.18 15.22 $\pm$ 0.11 16.51 $\pm$ 0.04 15.28 $\pm$ 0.21 0.73 $\pm$ 0.0 15.96 $\pm$ 0.0
Salman ResNet18 14.15 $\pm$ 0.22 2.04 $\pm$ 0.19 11.95 $\pm$ 0.00 10.58 $\pm$ 0.14 12.54 $\pm$ 0.02 12.39 $\pm$ 0.04 14.4 $\pm$ 0.17 12.34 $\pm$ 0.05 13.83 $\pm$ 0.14 0.79 $\pm$ 0.0 12.44 $\pm$ 0.0
Liu ConvNeXt-L 9.23 $\pm$ 0.13 0.42 $\pm$ 0.12 12.02 $\pm$ 0.00 4.94 $\pm$ 0.06 9.71 $\pm$ 0.01 7.28 $\pm$ 0.04 9.1 $\pm$ 0.06 9.46 $\pm$ 0.00 9.74 $\pm$ 0.0 0.36 $\pm$ 0.0 9.67 $\pm$ 0.0
Liu ConvNeXtV2-L+Swin-L 6.69 $\pm$ 0.08 0.88 $\pm$ 0.15 11.77 $\pm$ 0.01 5.33 $\pm$ 0.11 7.54 $\pm$ 0.03 5.09 $\pm$ 0.04 7.0 $\pm$ 0.02 7.57 $\pm$ 0.00 7.07 $\pm$ 0.09 0.37 $\pm$ 0.0 7.59 $\pm$ 0.0
Liu Swin-L 8.16 $\pm$ 0.09 0.55 $\pm$ 0.13 12.55 $\pm$ 0.00 4.4 $\pm$ 0.08 8.95 $\pm$ 0.01 6.78 $\pm$ 0.07 8.21 $\pm$ 0.01 8.89 $\pm$ 0.02 8.72 $\pm$ 0.04 0.31 $\pm$ 0.0 8.94 $\pm$ 0.0

### v2

@@ B-C Reproduction Package and Availability
TABLE XXII: Success rates of black-box attacks using vanilla surrogates against a vanilla target and robust targets.
Standard ResNet50 51.62 $\pm$ 0.0 65.95 $\pm$ 0.75 41.5 $\pm$ 0.35 61.22 $\pm$ 0.18 61.10 $\pm$ 0.13 64.75 $\pm$ 0.17 7.11 $\pm$ 0.25 49.36 $\pm$ 0.14 87.39 $\pm$ 0.26 81.08 $\pm$ 0.49 89.56 $\pm$ 0.09
Salman ResNet18 0.98 $\pm$ 0.0 1.51 $\pm$ 0.03 2.04 $\pm$ 0.06 1.29 $\pm$ 0.05 1.35 $\pm$ 0.06 1.88 $\pm$ 0.08 0.83 $\pm$ 0.03 1.13 $\pm$ 0.08 2.17 $\pm$ 0.05 2.36 $\pm$ 0.04 4.06 $\pm$ 0.06
Liu ConvNeXt-L 0.6 $\pm$ 0.0 0.92 $\pm$ 0.05 1.19 $\pm$ 0.1 0.7 $\pm$ 0.02 0.8 $\pm$ 0.04 0.99 $\pm$ 0.03 0.18 $\pm$ 0.04 0.48 $\pm$ 0.09 1.05 $\pm$ 0.02 1.16 $\pm$ 0.0 1.47 $\pm$ 0.01
Liu ConvNeXtV2-L+Swin-L 2.0 $\pm$ 0.0 2.79 $\pm$ 0.07 2.08 $\pm$ 0.04 2.51 $\pm$ 0.02 2.45 $\pm$ 0.05 2.52 $\pm$ 0.06 0.63 $\pm$ 0.03 1.7 $\pm$ 0.19 2.57 $\pm$ 0.07 4.77 $\pm$ 0.13 2.94 $\pm$ 0.12
Liu Swin-L 0.51 $\pm$ 0.0 0.88 $\pm$ 0.03 1.22 $\pm$ 0.05 0.74 $\pm$ 0.02 0.74 $\pm$ 0.04 0.93 $\pm$ 0.05 0.18 $\pm$ 0.0 0.53 $\pm$ 0.03 1.1 $\pm$ 0.11 1.26 $\pm$ 0.02 1.25 $\pm$ 0.01
@@
TABLE XXIII: Success rates of black-box attacks using robust surrogates against a vanilla target and robust targets.
Standard ResNet50 15.96 $\pm$ 0.0 15.28 $\pm$ 0.21 8.48 $\pm$ 0.18 16.47 $\pm$ 0.02 16.51 $\pm$ 0.04 15.22 $\pm$ 0.11 0.73 $\pm$ 0.0 17.95 $\pm$ 0.36 81.34 $\pm$ 0.43 17.94 $\pm$ 0.00 67.54 $\pm$ 0.13
Salman ResNet18 12.44 $\pm$ 0.0 13.83 $\pm$ 0.14 12.39 $\pm$ 0.04 12.54 $\pm$ 0.02 12.34 $\pm$ 0.05 14.4 $\pm$ 0.17 0.79 $\pm$ 0.0 14.15 $\pm$ 0.22 2.78 $\pm$ 0.21 11.95 $\pm$ 0.00 10.58 $\pm$ 0.14
Liu ConvNeXt-L 9.67 $\pm$ 0.0 9.74 $\pm$ 0.0 7.28 $\pm$ 0.04 9.71 $\pm$ 0.01 9.46 $\pm$ 0.00 9.1 $\pm$ 0.06 0.36 $\pm$ 0.0 9.23 $\pm$ 0.13 1.6 $\pm$ 0.11 12.02 $\pm$ 0.00 4.94 $\pm$ 0.06
Liu ConvNeXtV2-L+Swin-L 7.59 $\pm$ 0.0 7.07 $\pm$ 0.09 5.09 $\pm$ 0.04 7.54 $\pm$ 0.03 7.57 $\pm$ 0.00 7.0 $\pm$ 0.02 0.37 $\pm$ 0.0 6.69 $\pm$ 0.08 4.9 $\pm$ 0.06 11.77 $\pm$ 0.01 5.33 $\pm$ 0.11
Liu Swin-L 8.94 $\pm$ 0.0 8.72 $\pm$ 0.04 6.78 $\pm$ 0.07 8.95 $\pm$ 0.01 8.89 $\pm$ 0.02 8.21 $\pm$ 0.01 0.31 $\pm$ 0.0 8.16 $\pm$ 0.09 1.57 $\pm$ 0.04 12.55 $\pm$ 0.00 4.4 $\pm$ 0.08